how do you show the equation of a line that passes through point (0, b)and (1,b + m)is y=mx+b and how do you explain that point (-1, b -m) also is on the line?
(linear equations in slope intercept from)

Answer by mananth(16949)

(0,b) (1,b+m)are the points
(b+m-b)/1 = m is the slope
The equation of line
(y-y1)= m(x-x1)
plug (0,b) & m in the equation
(y-b)=m(x-0)
y-b=mx
y=mx+b
---
substitute -1, b-m in the equation y= mx+b
b-m = m*-1+b
b-m = b-m
so the point is on the line